OpenStudy (anonymous):

Please explain how to do this it is greek to me and need help find an equivalent expression with the given denominator 6/q=?/q to the power of 2

OpenStudy (anonymous):

multiply top and bottom by q to get \[\frac{6}{q}\times \frac{q}{q}=\frac{6q}{q^2}\]

OpenStudy (anonymous):

like writing \[\frac{6}{5}=\frac{30}{25}\]by multiplying top and bottom by 5

OpenStudy (anonymous):

so 6/q is 30/25